A quality control engineer at a potato chip company tests the bag filling machine by weighing bags of potato chips. Not every bag contains exactly the same weight. But if more than 10% of bags are over-filled then they stop production to fix the machine. They define over-filled to be more than 1 ounce above the weight on the package.

The engineer weighs 94 bags and finds that 18 of them are over-filled. He plans to test the hypotheses H0: p = 0.1 versus Ha: p > 0.1. What is the test statistic?
Answer #1

\hat p = x / n = 18 /94=0.191

P0 = 0.10

1 - P0 = 1 - 0.10=0.90

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.191 - 0.10/ [\sqrt(0.10*0.90) /94 ] 0.03094

= 2.9449

z=2.94